OpenStudy (anonymous):

(2x^4y)/(6xy^2) Help!

OpenStudy (nubeer):

\[x ^{2}\times x ^{3 } = x ^{2+3}\] if being divide \[\frac{ x ^{3} }{ x ^{2} } = x ^{3-2}\]

OpenStudy (nubeer):

\[\frac{ 2x ^{4}y }{ 6xy ^{2}}\] now apply the same technique on your question.

OpenStudy (anonymous):

I know how to do that for ex. if the question was \[6x^4y^2\div2xy \] it would be 3x^3y I just get confused when I have numbers and exponents on the top that are greater than the ones on the bottom.

OpenStudy (anonymous):

**less than the ones on the bottom

OpenStudy (nubeer):

doesn't matter .. if you have a power in numerator less then denominator you can write it as y^2/y^3 \[\frac{ 1 }{ y } \] OR \[ y ^{-1}\]

OpenStudy (nubeer):

both of these are same things.. so u can write them either way both are correct.

OpenStudy (anonymous):

Would \[\frac{ 1x^3 }{ 3y } \] Would that be correct?

OpenStudy (nubeer):

yes..